OpenStudy (amour.anohs):

Rearrange the formula c2 = a2 + b2 for b.

OpenStudy (amour.anohs):

c2 = a2 + b2 c2 - a2 = b2 sqrt {c2 - a2} = b b =sqrt{ c2 - a2} is correct.

OpenStudy (anonymous):

yep

candycove (candycove):

omg